# Systematic effects in galaxy–galaxy lensing with DESI

### v1

@@ 1 Introduction
Gravitational lensing by foreground galaxies induces shear distortions in the apparent images of photometric galaxies behind it, giving rise to a non-zero mean tangential shear $\gamma_{\mathrm{t}}$. While this gravitational effect is most often the primary contributor to the apparent shape distortions around foreground spectroscopic galaxies, various systematic effects can also influence galaxy–galaxy lensing measurements. These effects include shear biases, magnification effects, incompleteness in the spectroscopic samples, photometric redshift dilution, and intrinsic alignment (IA). IA describes the intrinsic correlations between the shapes of nearby galaxies. As such, IA can mimic gravitational lensing and is one of the most significant sources of systematic errors in weak lensing. It has been studied extensively, particularly in the context of photometric surveys measuring the so-called cosmic shear, i.e., shape correlations of nearby galaxies. For an overview of IA see Lamman et al. (2024) and for comprehensive reviews see (Joachimi et al., 2015; Kiessling et al., 2015; Troxel & Ishak, 2015). We note that another important effect in the interpretation of gravitational lensing is baryonic feedback (see, e.g. Leauthaud et al., 2017; Lange et al., 2019; Amodeo et al., 2021). However, unlike the aforementioned effects, baryonic feedback alters the intrinsic matter field around the foreground galaxies. Thus, it is typically addressed by including baryonic physics in the modeling framework and will not be explored in the present paper which is concerned with potential biases in the measurement of lensing.
@@
Systematic effects in galaxy–galaxy lensing measurements have been investigated in the past using theoretical calculations, simulations and observational data (see, e.g., Mandelbaum et al., 2005, 2013; Miyatake et al., 2015; Simet & Mandelbaum, 2015; Leauthaud et al., 2017; Speagle et al., 2019; Unruh et al., 2020; Prat et al., 2022; Makiya & Sunayama, 2022; Yao et al., 2023b). However, given the high-precision measurements expected via a joint analysis of DESI, DES, HSC, and KiDS, further investigation is warranted. Our work extends previous studies by utilizing sophisticated, high-resolution simulations, the Buzzard mocks (DeRose et al., 2019), that realistically simulate the above systematic effects. Additionally, many of the systematic effects discussed above depend on the data sets in question, such as the precision of photometric redshifts for the weak lensing survey or the nature and magnitude of incompleteness for spectroscopic samples. We are thus particularly interested in quantifying the importance of these systematic effects for combined studies with DESI, especially in the context of the upcoming DESI galaxy–galaxy lensing analysis. In addition to estimating the magnitude of these systematic effects, we will also investigate possible mitigation strategies. Given the statistical constraining power of cross-correlation studies of DESI with DES, HSC, and KiDS (Yuan et al., 2024), we expect that the systematic error will need to be at the level of, at most, a few percent for cosmological studies to be statistically limited (Yao et al., 2023a). We also refer the reader to a companion study, Yuan et al. (2024), that focuses on another systematic, redshift evolution effects, in combined studies of DESI clustering and gravitational lensing. Finally, we note that detailed image simulations needed for characterizing shear biases and the effects of blending are beyond the scope of this work. Instead, those effects have been studied by the DES, HSC, and KiDS collaborations in their respective studies (Mandelbaum et al., 2018; Giblin et al., 2021; MacCrann et al., 2022).
@@ 4.5 Fiber assignment
The bias in the selection of spectroscopic targets can be resolved by weighing each lens galaxy by its individual inverse probability (IIP) of being assigned a fiber, $w_{\mathrm{sys}}\rightarrow w_{\mathrm{sys}}\,w_{\mathrm{IIP}}$. The IIP can be estimated from Monte-Carlo simulations of the DESI fiber assignment pipeline on the photometric data and is one over the fraction of realizations in which each target is assigned a fiber. The idea of IIP-weighting is analogous to weighting by the pairwise-inverse-probability (PIP) weights for galaxy clustering (Bianchi et al., 2018). The reason for using the individual instead of the pairwise probability is that in galaxy–galaxy lensing, only the lens is assigned a fiber, whereas for clustering both targets are subject to fiber incompleteness. In Fig. 7, we show that this IIP-weighting resolves any biases arising from fiber assignments: within the uncertainties the lensing signal around all targets and only the subset of targets with spectroscopic redshifts is identical to within $\lesssim 1\%$ once the IIP-weights are applied. Finally, we also confirmed that this correction gives unbiased estimates even if the DESI spectroscopic coverage was very incomplete, i.e., only covering the sky with one pass.
@@ 4.6 Lens magnification
$\Delta\Sigma_{\mathrm{lm}}(r_{\mathrm{p}})=$ $2\Sigma_{\mathrm{c}}(\overline{z_{\mathrm{l}}},\overline{z_{\mathrm{s}}})(\alpha_{\mathrm{l}}-1)$ (28)
$\gamma_{\mathrm{LSS}}\left(\frac{r_{\mathrm{p}}}{D_{\mathrm{C}}(\overline{z_{\mathrm{l}}})},\overline{z_{\mathrm{l}}},\overline{z_{\mathrm{s}}}\right)\,,$
@@ 5 Discussion
In Fig. 13, we show the expected significance of the most relevant systematic effects as a function of the minimum scale analyzed. We find that the effect of IA is most relevant if one measures $\gamma_{\mathrm{t}}$ for all lens-source bin combinations. This may include bin combinations where sources and lenses significantly overlap in redshift. In this case, the mean tangential shear may be dominated by IA. However, in the same figure, we also show that the impact of IA can be reduced dramatically if one employs suitable lens-source cuts, as done for the mock $\Delta\Sigma$ measurements. In this case, the impact of IA can be reduced to $\Delta\chi^{2}\sim 1-3$, although it may still be significant. The residual impact is the strongest for DES owing to their relatively imprecise redshift estimates which leads to very wide tomographic bins. We recommend that all studies, particularly those not employing lens-source redshift cuts, should include a model for IA in their analysis. Finally, we note that our estimate of the IA contamination is based on the current observational constraints on IA via the Non-Linear Alignment model (NLA; Bridle & King, 2007). However, more complex models such as the (TATT, Blazek et al., 2019) or halo-based models (Fortuna et al., 2021; Van Alfen et al., 2023), may describe the behavior of IA more accurately, especially on small-scale. Furthermore, the strength of IA is poorly constrained empirically. While our NLA amplitude, $A_{\textrm{IA}}=0.5$, is consistent with current data, $A_{\textrm{IA}}=1.0$ is not strongly excluded. Thus, the impact of IA on galaxy–galaxy lensing may realistically be much higher than our mock results indicate. This further strengthens the conclusion that IA-contamination needs to be considered for galaxy–galaxy lensing estimates.
@@ 6 Conclusion
This work performs analysis choices for several upcoming studies cross-correlating DESI targets with weak lensing source catalogs. Among others, we plan to perform a unified analysis of cosmic shear, galaxy–galaxy lensing, and galaxy clustering, a so-called $3\times 2$pt-analysis, with DESI, DES, HSC, and KiDS. Such an analysis is expected to result in leading constraints on cosmic structure growth and shed light on the so-called $S_{8}$-tension (Abdalla et al., 2022). Similarly, we will perform measurements of $\Delta\Sigma$ around common sets of DESI lenses. Since the measured $\Delta\Sigma$ amplitude has no dependence on the source catalog properties in the absence of the systematic effects discussed here, comparing measured $\Delta\Sigma$ amplitudes can be used as a powerful tool to empirically test the robustness of weak lensing measurements (Leauthaud et al., 2022). The work presented here is a first step towards such future studies with DESI data.
@@ Acknowledgments
We acknowledge the use of the lux supercomputer at UC Santa Cruz, funded by NSF MRI grant AST 1828315. We acknowledge support from the Leinweber Center for Theoretical Physics, NASA grant under contract 19-ATP19-0058, and DOE under contract DE-FG02-95ER40899.

### v2

@@ 1 Introduction
Gravitational lensing by foreground galaxies induces shear distortions in the apparent images of photometric galaxies behind it, giving rise to a non-zero mean tangential shear $\gamma_{\mathrm{t}}$. While this gravitational effect is most often the primary contributor to the apparent shape distortions around foreground spectroscopic galaxies, various systematic effects can also influence galaxy–galaxy lensing measurements. These effects include shear biases, magnification effects, incompleteness in the spectroscopic samples, photometric redshift dilution, and intrinsic alignment (IA). IA describes the intrinsic correlations between the shapes of nearby galaxies. As such, IA can mimic gravitational lensing and is one of the most significant sources of systematic errors in weak lensing. It has been studied extensively, particularly in the context of photometric surveys measuring the so-called cosmic shear, i.e., shape correlations of nearby galaxies. For an overview of IA see Lamman et al. (2024) and for comprehensive reviews see (Joachimi et al., 2015; Kiessling et al., 2015; Troxel & Ishak, 2015). We note that another important effect in the interpretation of gravitational lensing is baryonic feedback (see, e.g., Leauthaud et al., 2017; Lange et al., 2019; Amodeo et al., 2021; Beltz-Mohrmann & Berlind, 2021; Amon et al., 2023). However, unlike the aforementioned effects, baryonic feedback alters the intrinsic matter field around the foreground galaxies. Thus, it is typically addressed by including baryonic physics in the modeling framework and will not be explored in the present paper which is concerned with potential biases in the measurement of lensing.
@@
Systematic effects in galaxy–galaxy lensing measurements have been investigated in the past using theoretical calculations, simulations and observational data (see, e.g., Mandelbaum et al., 2005, 2013; Miyatake et al., 2015; Simet & Mandelbaum, 2015; Leauthaud et al., 2017; Speagle et al., 2019; Unruh et al., 2020; Prat et al., 2022; Makiya & Sunayama, 2022; Yao et al., 2023b; Amon et al., 2023). However, given the high-precision measurements expected via a joint analysis of DESI, DES, HSC, and KiDS, further investigation is warranted. Our work extends previous studies by utilizing sophisticated, high-resolution simulations, the Buzzard mocks (DeRose et al., 2019), that realistically simulate the above systematic effects. Additionally, many of the systematic effects discussed above depend on the data sets in question, such as the precision of photometric redshifts for the weak lensing survey or the nature and magnitude of incompleteness for spectroscopic samples. We are thus particularly interested in quantifying the importance of these systematic effects for combined studies with DESI, especially in the context of the upcoming DESI galaxy–galaxy lensing analysis. In addition to estimating the magnitude of these systematic effects, we will also investigate possible mitigation strategies. Given the statistical constraining power of cross-correlation studies of DESI with DES, HSC, and KiDS (Yuan et al., 2024), we expect that the systematic error will need to be at the level of, at most, a few percent for cosmological studies to be statistically limited (Yao et al., 2023a). We also refer the reader to a companion study, Yuan et al. (2024), that focuses on another systematic, redshift evolution effects, in combined studies of DESI clustering and gravitational lensing. Finally, we note that detailed image simulations needed for characterizing shear biases and the effects of blending are beyond the scope of this work. Instead, those effects have been studied by the DES, HSC, and KiDS collaborations in their respective studies (Mandelbaum et al., 2018; Giblin et al., 2021; MacCrann et al., 2022).
@@ 4.5 Fiber assignment
The bias in the selection of spectroscopic targets can be resolved by weighing each lens galaxy by its individual inverse probability (IIP) of being assigned a fiber, $w_{\mathrm{sys}}\rightarrow w_{\mathrm{sys}}\,w_{\mathrm{IIP}}$. The IIP can be estimated from Monte-Carlo simulations of the DESI fiber assignment pipeline on the photometric data and is one over the fraction of realizations in which each target is assigned a fiber. The idea of IIP-weighting is analogous to weighting by the pairwise-inverse-probability (PIP) weights for galaxy clustering (Bianchi et al., 2018). The reason for using the individual instead of the pairwise probability is that in galaxy–galaxy lensing, only the lens is assigned a fiber, whereas for clustering both targets are subject to fiber incompleteness. In Fig. 7, we show that this IIP-weighting resolves any biases arising from fiber assignments: within the uncertainties the lensing signal around all targets and only the subset of targets with spectroscopic redshifts is identical to within $\lesssim 1\%$ once the IIP-weights are applied. Finally, we also confirmed that this correction gives unbiased estimates even if the DESI spectroscopic coverage was very incomplete, i.e., only covering the sky with one pass. The results presented here are fully consistent with earlier studies by Makiya & Sunayama (2022) on the impact of fiber incompleteness on the galaxy–galaxy lensing amplitude around galaxies observed with the Prime Focus Spectrograph (PFS).
@@ 4.6 Lens magnification
$\Delta\Sigma_{\mathrm{lm}}(r_{\mathrm{p}})=2(\alpha_{\mathrm{l}}-1)\,\overline{\Sigma_{\mathrm{c}}}\,\gamma_{\mathrm{LSS}}\left(\frac{r_{\mathrm{p}}}{D_{\mathrm{C}}(\overline{z_{\mathrm{l}}})},\overline{z_{\mathrm{l}}},\overline{z_{\mathrm{s}}}\right)\,,$ (28)
@@
Similarly, $\overline{\Sigma_{\mathrm{c}}}$ is the mean critical surface density.
@@ 5 Discussion
In Fig. 13, we show the expected significance of the most relevant systematic effects as a function of the minimum scale analyzed. We find that the effect of IA is most relevant if one measures $\gamma_{\mathrm{t}}$ for all lens-source bin combinations. This may include bin combinations where sources and lenses significantly overlap in redshift. In this case, the mean tangential shear may be dominated by IA. However, in the same figure, we also show that the impact of IA can be reduced dramatically if one employs suitable lens-source cuts, as done for the mock $\Delta\Sigma$ measurements. In this case, the impact of IA can be reduced to $\Delta\chi^{2}\sim 1-3$, although it may still be significant. The residual impact is the strongest for DES owing to their relatively imprecise redshift estimates which leads to very wide tomographic bins. We recommend that all studies, particularly those not employing lens-source redshift cuts, should include a model for IA in their analysis. Finally, we note that our estimate of the IA contamination is based on the current observational constraints on IA via the Non-Linear Alignment model (NLA; Bridle & King, 2007). However, more complex models such as the (TATT, Blazek et al., 2019) or halo-based models (Fortuna et al., 2021; Van Alfen et al., 2023), may describe the behavior of IA more accurately, especially on small-scale. Furthermore, the strength of IA is poorly constrained empirically. While our NLA amplitude, $A_{\textrm{IA}}=0.5$, is consistent with current data, $A_{\textrm{IA}}=1.0$ is not strongly excluded. Finally, it has been found in simulations that the IA amplitude increases with stellar mass and redshift (see Delgado et al., 2023, and references therein). As a result, flux-limited lensing surveys may show stronger IA amplitudes at high-redshift (see, e.g., Dark Energy Survey and Kilo-Degree Survey Collaboration et al., 2023), leading to a further underestimation of the IA contamination, especially around LRGs. Ultimately, the impact of IA on galaxy–galaxy lensing may realistically be much higher than our mock results indicate. This further strengthens the conclusion that IA-contamination needs to be considered for galaxy–galaxy lensing estimates.
@@ 6 Conclusion
This work performs analysis choices for several upcoming studies cross-correlating DESI targets with weak lensing source catalogs. Among others, we plan to perform a unified analysis of cosmic shear, galaxy–galaxy lensing, and galaxy clustering, a so-called $3\times 2$pt-analysis, with DESI, DES, HSC, and KiDS. Such an analysis is expected to result in leading constraints on cosmic structure growth and shed light on the so-called $S_{8}$-tension (Abdalla et al., 2022). Similarly, we will perform measurements of $\Delta\Sigma$ around common sets of DESI lenses. Since the measured $\Delta\Sigma$ amplitude has no dependence on the source catalog properties in the absence of the systematic effects discussed here, comparing measured $\Delta\Sigma$ amplitudes can be used as a powerful tool to empirically test the robustness of weak lensing measurements (see Leauthaud et al., 2022; Amon et al., 2023, for an application to BOSS). The work presented here is a first step towards such future studies with DESI data.
@@ Acknowledgments
We thank Ji Yao, Alexandra Amon, Tomomi Sunayama, and the referee for insightful comments that improved this paper. We acknowledge the use of the lux supercomputer at UC Santa Cruz, funded by NSF MRI grant AST 1828315. We acknowledge support from the Leinweber Center for Theoretical Physics, NASA grant under contract 19-ATP19-0058, and DOE under contract DE-FG02-95ER40899.